A food contains 3 g of total fat and 1 g of saturated fat. How many kcalories of fat does the food contain?

  

27

   

12

   

16

   

36

   

4
Answer #1

Answer is 36 kcal (option 4)

1 gm of total fat will provide 9kcal of energy.

1 gm of saturated fat will also give 9kcal energy.

So 3gm

3gm of total fat will give (3*9)=27 kcal energy.

1gm of saturated fat will give (1*9)= 9kcal energy.

Total energy = (27+9)= 36 kcal energy.

So,option a,b,c and e are not correct.

As,it is 36kcal energy.

Hence, option d is correct.